Ch 02: Using Financial Statements and Budgets They will put the $5,000 in an account along with money they will deposit annually. They dont know how much that annual deposit should be, so theyve asked you to calculate it They have found a savings institution that will pay 6% interest. The Colbys have set a goal to have $25,000 for a down payment on a house in 5 years. Use the scenarios along with the following factor table data to answer each of the questions. Note that the complete Future Value and Future Value Annulty tables (as well as the Present Value and Present Valu e Annuity tables) are located in the appendlx in your text.
Ch 02: Assignment-Using Financial Statements and Budgets 596 6% 0% 1.050 1.060 1.08 1.102 1.120 1.166 1.158 1.190 1.260o 1.216 1.260 1.360 1.276 1.340 1.469 1.340 1.420 1.587 1.477 1.590 1.851 1.629 1.790 2.159 10 5% 6% 8% 1.000 1.000 1.000 3.152 3.180 3.246 4.310 4.380 4.506 5.526 6.802 6.970 7.336 9.549 9.890 10.637 12.578 13.180 14.487 630 5.867 10 What is the amount of money the Hardys will need to deposit annually (rounded to the nearest two decimal places) to achieve their down-payment goal? What is the amount of money the Colbys will need to deposit annually (rounded to the nearest two decimal places) to achieve their down-payment goal?
Answer #1
HARDYS FAMILY:
FV of 5000 after 5 years = 5000*1.34 = $             6,700
Balance required at EOY 5, 25000-6700 = $          18,300
Annual deposit required = 18300/5.63 = $             3,250
COLBYS:
Annual deposit required = 25000/5.63 = $             4,440